# Congestion-Approximators from the Bottom Up

### v1

@@ 2.1 Bottom-Up Construction
As mentioned previously, the partition $\mathcal{P}_{2}$ is a weak-expander decomposition of the graph, so it can be computed in nearly-linear time using off-the-shelf expander decomposition algorithm that avoid any black-box call to max-flow [SW19]. For partitions $\mathcal{P}_{3}$ onwards, expansion is measured with respect to a subset of edges, so simple expander decomposition algorithms no longer suffice. Instead, we have to resort to expander decomposition algorithms that make black-box calls to (approximate) max-flow. Naïvely, these max-flows can be computed recursively, resulting in a congestion approximator algorithm that makes recursive calls to max-flow, similar to [RST14]. Our key insight is that these max-flow instances are actually well-structured enough that recursion is unnecessary. In particular, to construct partition $\mathcal{P}_{i+1}$, the first $i$ partitions $\mathcal{P}_{1},\mathcal{P}_{2},\ldots,\mathcal{P}_{i}$—which the algorithm has already computed—can be converted to a pseudo-congestion approximator, which then translates to a max-flow algorithm sufficient for these well-structured instances.
@@ 3 Preliminaries
Given a graph $H=(U,F)$, a congestion-approximator $\mathcal{C}$ of quality $\alpha$ is a collection of subsets of $U$ such that for any demand $\mathbf{b}$ satisfying $|\mathbf{b}(C)|\leq\delta_{H}C$ for all $C\in\mathcal{C}$, there is a flow routing demand $\mathbf{b}$ with congestion $\alpha$. Through Sherman’s framework [She13, She17], a congestion-approximator of quality $\alpha$ translates to a $(1+\epsilon)$-approximate maximum flow algorithm with running time $\tilde{O}(\epsilon^{-1}\alpha m)$. [^12] The $\tilde{O}(\cdot)$ notation hides polylogarithmic factors in $n$. In our paper, it is most convenient to work with a stronger variant of approximate min-cut/max-flow called fair cut/flow [LNPS23], which is formally defined in Section 5.2.
@@ 4 Bottom-Up Congestion-Approximator
In this section, we formally state our congestion-approximator construction and prove its quality guarantee. See Figure 1 for an illustration. The most important distinction is that we define and analyze mixing on vertex weightings, not edges. This simplifies the notation since we can avoid working with the subdivision graph in [RST14]. For example, in property (2) below, the mixing of the vertex weighting $\mathrm{deg}_{\partial\mathcal{P}_{i}\cup\partial C}|_{C}$ is conceptually equivalent to the mixing of the edges of $\partial\mathcal{P}_{i}$ internal to $C$ together with the boundary edges $\partial C$.
@@ 5 Partitioning Algorithm
At a high level, our algorithm proceeds similarly to expander decomposition algorithms like [SW19], where expanders are defined with respect to the vertex weighting $\deg_{\partial\mathcal{P}_{L}}$ of partition $\mathcal{P}_{L}$. We iteratively decompose the graph using the cut-matching game of [KRV09a]: for each cluster of the decomposition, we either compute a low-conductance cut or certify that the current cluster is mixing (or in [SW19] terms, a nearly expander). The matching step of the cut-matching game requires a call to approximate min-cut/max-flow, but recall that the partitions $\mathcal{P}_{1},\ldots,\mathcal{P}_{L}$ only form a pseudo-congestion-approximator. Luckily, we can modify the graph in a way that the pseudo-congestion-approximator can be adapted to an actual congestion-approximator for the new graph. We then black-box a cut/flow algorithm on this new graph, and then modify the cut and flow to fit the old graph in a way that suffices for the cut-matching game.
@@
    In Section 5.2, we cite the (approximate) fair cut/flow algorithm of [LNPS23], which computes a cut/flow pair with desirable properties given a congestion-approximator.
@@
    In Section 5.3, we introduce the cut-matching game as well as a trimming procedure similar to [SW19]. Both the cut-matching game and the trimming step use the fair cut/flow algorithm (Section 5.2) on the modified graphs for which we have a congestion-approximator (Section 5.1).
@@ 5.2 Fair Cut/Flow Algorithm
Given a congestion-approximator, the most convenient min-cut/max-flow algorithm is the fair cut/flow algorithm of [LNPS23]. We state the definition of a fair cut/flow and then cite the main result of [LNPS23].
@@
Theorem 5.9 (Fair cut/flow algorithm [LNPS23]).
@@ 5.3 Cut-Matching Game and Trimming
We follow the cut-matching game treatment in [SW19]: either find a “balanced” cut of small capacity, or ensure that a “large” part of the graph mixes with low congestion. The following lemma is similar to Theorem 2.2 of [SW19] with one key difference: there is no built-in flow subroutine, so the algorithm makes black-box calls to the fair cut/flow algorithm of Theorem 5.10.
In past work [RST14, AKL+24], the analysis of the cut-matching game for capacitated graphs has only been sketched, referencing the fact that a capacitated graph can be modelled by an uncapacitated graph with parallel edges (at a cost). For completeness, we provide a full proof of this capacitated case in Appendix A.
@@
Property (3) asserts that either an approximately balanced cut is found, or the vertex weighting $\mathbf{d}|_{A\setminus R}$ mixes with low congestion (with high probability). In our algorithm for Theorem 5.1, we actually want the weighting $\mathbf{d}_{A\setminus R}+\deg_{\partial R}$ to mix in the second case. To guarantee this stronger property, we augment the set $R$ into $R\cup B$ through one additional call to the fair cut/flow algorithm of Theorem 5.10. The algorithm is similar to the flow-based expander trimming procedure in [SW19]. For completeness, we defer the algorithm and proof to Appendix B. Note that the setting, including assumption 1, is the same as Theorem 5.11.
@@ 6 Approximate Maximum Flow
From Theorem 5.1 and Claim 5.2, we obtain an algorithm that constructs a congestion-approximator of quality $O(\log^{10}(nW))$ in $\tilde{O}(m)$ time. Recall that Sherman’s framework [She13, She17] translates a congestion-approximator of quality $\alpha$ to a $(1+\epsilon)$-approximate maximum flow algorithm with running time $\tilde{O}(\epsilon^{-1}\alpha m)$. Thus, for any parameter $\epsilon>0$, we obtain a $(1+\epsilon)$-approximate maximum flow algorithm with running time $\tilde{O}(\epsilon^{-1}m)$.
@@ Appendix A Cut-Matching Game
Our setup resembles Appendix B of [SW19] with a few minor changes. In particular, since our flow routine is more restrictive, we have to adapt the algorithm to handle our flow outputs.
We begin with notation from [SW19]. For simplicity, we avoid working with the subdivision graph in [RST14, SW19]. Define a $A$-commodity flow as a multi-commodity flow where each vertex $v\in A$ is the source of quantity $\mathbf{d}(v)$ of its distinct flow commodity. Only for analysis, we consider a $A\times A$ flow-matrix $\mathbf{F}\in\mathbb{R}^{A\times A}_{\geq 0}$ which encodes information about a $A$-commodity flow. We say that $\mathbf{F}$ is routable with congestion $c$ if there exists a $A$-commodity flow $f$ such that, simultaneously for all $u,v\in A$, we have that $u$ can send quantity $\mathbf{F}(u,v)$ of its own commodity to $v$, and the amount of flow through each edge is at most $c$.
@@ A.2 Computing the Set and Flow
In this subsection, we describe a round of the algorithm in detail, following Lemma B.3 of [SW19] with some minor changes. Recall that on each iteration $t\in[T]$, the algorithm first computes a partition of $A_{t-1}$ into $A_{t}^{\ell},A_{t}^{r}$, and then computes a set $S_{t}$ and flow $f_{t}$ satisfying certain properties. The algorithm then updates $A_{t}\leftarrow A_{t-1}\setminus S_{t}$ and $R_{t}\leftarrow R_{t-1}\cup S_{t}$. The choice of partition $A_{t}^{\ell},A_{t}^{r}$ will depend on an implicitly represented flow-matrix $\mathbf{F}_{t}$ that is useful for the analysis.
@@ A.2.1 Constructing the partition
We conclude this section with a lemma about the behavior of projections onto a random unit vector. Since the lemma is standard in high-dimensional geometry, we omit the proof and refer to Lemma 3.4 of [KRV09a].
Lemma A.6 (Lemma 3.4 of [KRV09a]).
@@ A.2.2 Max-flow/min-cut call
The algorithm defines the flow $f_{t}$ as follows. First, scale the flow $f$ by factor $1/(12\phi)$ and restrict the flow to edges in $G[A\setminus S_{t}]$. In other words, the flow on edges incident to $S_{t}\cup\{s,t,x\}$ are removed. Then, decompose the flow $f_{t}$ into paths (using, for example, a dynamic tree [ST81]). For each vertex $v\in A_{t}^{\ell}\setminus S_{t}$, remove enough paths starting at $v$ until it is the start of exactly $\mathbf{d}(v)/12$ total capacity of paths, and for each vertex $v\notin A_{t}^{\ell}\setminus S_{t}$, remove all paths starting at $v$; let the remaining flow be $f_{t}$. The claim below shows that this last step is always possible, and that $f_{t}$ satisfies the condition below as required by step ii.
@@ A.3 Analyzing the Potential Decrease
The following lemma adapts Lemma 3.4 of [RST14] to the capacitated case.
@@ Appendix B Expander Trimming
Our setup closely resembles Section 3.2 of [SW19], except that we replace the exact min-cut/max-flow oracle in their “slow-trimming” procedure with an approximate one. We also avoid defining expanders and work exclusively with flow mixing, which simplifies the analysis.

### v2

@@ 2.1 Bottom-Up Construction
As mentioned previously, the partition $\mathcal{P}_{2}$ is a weak-expander decomposition of the graph, so it can be computed in nearly-linear time using off-the-shelf expander decomposition algorithm that avoid any black-box call to max-flow [34]. For partitions $\mathcal{P}_{3}$ onwards, expansion is measured with respect to a subset of edges, so simple expander decomposition algorithms no longer suffice. Instead, we have to resort to expander decomposition algorithms that make black-box calls to (approximate) max-flow. Naïvely, these max-flows can be computed recursively, resulting in a congestion approximator algorithm that makes recursive calls to max-flow, similar to [33]. Our key insight is that these max-flow instances are actually well-structured enough that recursion is unnecessary. In particular, to construct partition $\mathcal{P}_{i+1}$, the first $i$ partitions $\mathcal{P}_{1},\mathcal{P}_{2},\ldots,\mathcal{P}_{i}$—which the algorithm has already computed—can be converted to a pseudo-congestion approximator, which then translates to a max-flow algorithm sufficient for these well-structured instances.
@@ 3 Preliminaries
Given a graph $H=(U,F)$, a congestion-approximator $\mathcal{C}$ of quality $\alpha$ is a collection of subsets of $U$ such that for any demand $\mathbf{b}$ satisfying $|\mathbf{b}(C)|\leq\delta_{H}C$ for all $C\in\mathcal{C}$, there is a flow routing demand $\mathbf{b}$ with congestion $\alpha$. Through Sherman’s framework [36, 37], a congestion-approximator of quality $\alpha$ translates to a $(1+\epsilon)$-approximate maximum flow algorithm with running time $\tilde{O}(\epsilon^{-1}\alpha m)$. [^12] The $\tilde{O}(\cdot)$ notation hides polylogarithmic factors in $n$. In our paper, it is most convenient to work with a stronger variant of approximate min-cut/max-flow called fair cut/flow [25], which is formally defined in Section 5.2.
@@ 4 Bottom-Up Congestion-Approximator
In this section, we formally state our congestion-approximator construction and prove its quality guarantee. See Figure 1 for an illustration. The most important distinction is that we define and analyze mixing on vertex weightings, not edges. This simplifies the notation since we can avoid working with the subdivision graph in [33]. For example, in property (2) below, the mixing of the vertex weighting $\mathrm{deg}_{\partial\mathcal{P}_{i}\cup\partial C}|_{C}$ is conceptually equivalent to the mixing of the edges of $\partial\mathcal{P}_{i}$ internal to $C$ together with the boundary edges $\partial C$.
@@ 5 Partitioning Algorithm
At a high level, our algorithm proceeds similarly to expander decomposition algorithms like [34], where expanders are defined with respect to the vertex weighting $\deg_{\partial\mathcal{P}_{L}}$ of partition $\mathcal{P}_{L}$. We iteratively decompose the graph using the cut-matching game of [18]: for each cluster of the decomposition, we either compute a low-conductance cut or certify that the current cluster is mixing (or in [34] terms, a nearly expander). The matching step of the cut-matching game requires a call to approximate min-cut/max-flow, but recall that the partitions $\mathcal{P}_{1},\ldots,\mathcal{P}_{L}$ only form a pseudo-congestion-approximator. Luckily, we can modify the graph in a way that the pseudo-congestion-approximator can be adapted to an actual congestion-approximator for the new graph. We then black-box a cut/flow algorithm on this new graph, and then modify the cut and flow to fit the old graph in a way that suffices for the cut-matching game.
@@
    In Section 5.2, we cite the (approximate) fair cut/flow algorithm of [25], which computes a cut/flow pair with desirable properties given a congestion-approximator.
@@
    In Section 5.3, we introduce the cut-matching game as well as a trimming procedure similar to [34]. Both the cut-matching game and the trimming step use the fair cut/flow algorithm (Section 5.2) on the modified graphs for which we have a congestion-approximator (Section 5.1).
@@ 5.2 Fair Cut/Flow Algorithm
Given a congestion-approximator, the most convenient min-cut/max-flow algorithm is the fair cut/flow algorithm of [25]. We state the definition of a fair cut/flow and then cite the main result of [25].
@@
Theorem 5.9 (Fair cut/flow algorithm [25]).
@@ 5.3 Cut-Matching Game and Trimming
We follow the cut-matching game treatment in [34]: either find a “balanced” cut of small capacity, or ensure that a “large” part of the graph mixes with low congestion. The following lemma is similar to Theorem 2.2 of [34] with one key difference: there is no built-in flow subroutine, so the algorithm makes black-box calls to the fair cut/flow algorithm of Theorem 5.10.
In past work [33, 1], the analysis of the cut-matching game for capacitated graphs has only been sketched, referencing the fact that a capacitated graph can be modelled by an uncapacitated graph with parallel edges (at a cost). For completeness, we provide a full proof of this capacitated case in Appendix A.
@@
Property (3) asserts that either an approximately balanced cut is found, or the vertex weighting $\mathbf{d}|_{A\setminus R}$ mixes with low congestion (with high probability). In our algorithm for Theorem 5.1, we actually want the weighting $\mathbf{d}_{A\setminus R}+\deg_{\partial R}$ to mix in the second case. To guarantee this stronger property, we augment the set $R$ into $R\cup B$ through one additional call to the fair cut/flow algorithm of Theorem 5.10. The algorithm is similar to the flow-based expander trimming procedure in [34]. For completeness, we defer the algorithm and proof to Appendix B. Note that the setting, including assumption 1, is the same as Theorem 5.11.
@@ 6 Approximate Maximum Flow
From Theorem 5.1 and Claim 5.2, we obtain an algorithm that constructs a congestion-approximator of quality $O(\log^{10}(nW))$ in $\tilde{O}(m)$ time. Recall that Sherman’s framework [36, 37] translates a congestion-approximator of quality $\alpha$ to a $(1+\epsilon)$-approximate maximum flow algorithm with running time $\tilde{O}(\epsilon^{-1}\alpha m)$. Thus, for any parameter $\epsilon>0$, we obtain a $(1+\epsilon)$-approximate maximum flow algorithm with running time $\tilde{O}(\epsilon^{-1}m)$.
@@ Appendix A Cut-Matching Game
Our setup resembles Appendix B of [34] with a few minor changes. In particular, since our flow routine is more restrictive, we have to adapt the algorithm to handle our flow outputs.
We begin with notation from [34]. For simplicity, we avoid working with the subdivision graph in [33, 34]. Define a $A$-commodity flow as a multi-commodity flow where each vertex $v\in A$ is the source of quantity $\mathbf{d}(v)$ of its distinct flow commodity. Only for analysis, we consider a $A\times A$ flow-matrix $\mathbf{F}\in\mathbb{R}^{A\times A}_{\geq 0}$ which encodes information about a $A$-commodity flow. We say that $\mathbf{F}$ is routable with congestion $c$ if there exists a $A$-commodity flow $f$ such that, simultaneously for all $u,v\in A$, we have that $u$ can send quantity $\mathbf{F}(u,v)$ of its own commodity to $v$, and the amount of flow through each edge is at most $c$.
@@ A.2 Computing the Set and Flow
In this subsection, we describe a round of the algorithm in detail, following Lemma B.3 of [34] with some minor changes. Recall that on each iteration $t\in[T]$, the algorithm first computes a partition of $A_{t-1}$ into $A_{t}^{\ell},A_{t}^{r}$, and then computes a set $S_{t}$ and flow $f_{t}$ satisfying certain properties. The algorithm then updates $A_{t}\leftarrow A_{t-1}\setminus S_{t}$ and $R_{t}\leftarrow R_{t-1}\cup S_{t}$. The choice of partition $A_{t}^{\ell},A_{t}^{r}$ will depend on an implicitly represented flow-matrix $\mathbf{F}_{t}$ that is useful for the analysis.
@@ A.2.1 Constructing the partition
We conclude this section with a lemma about the behavior of projections onto a random unit vector. Since the lemma is standard in high-dimensional geometry, we omit the proof and refer to Lemma 3.4 of [18].
Lemma A.6 (Lemma 3.4 of [18]).
@@ A.2.2 Max-flow/min-cut call
The algorithm defines the flow $f_{t}$ as follows. First, scale the flow $f$ by factor $1/(12\phi)$ and restrict the flow to edges in $G[A\setminus S_{t}]$. In other words, the flow on edges incident to $S_{t}\cup\{s,t,x\}$ are removed. Then, decompose the flow $f_{t}$ into paths (using, for example, a dynamic tree [ST83]). For each vertex $v\in A_{t}^{\ell}\setminus S_{t}$, remove enough paths starting at $v$ until it is the start of exactly $\mathbf{d}(v)/12$ total capacity of paths, and for each vertex $v\notin A_{t}^{\ell}\setminus S_{t}$, remove all paths starting at $v$; let the remaining flow be $f_{t}$. The claim below shows that this last step is always possible, and that $f_{t}$ satisfies the condition below as required by step ii.
@@ A.3 Analyzing the Potential Decrease
The following lemma adapts Lemma 3.4 of [33] to the capacitated case.
@@ Appendix B Expander Trimming
Our setup closely resembles Section 3.2 of [34], except that we replace the exact min-cut/max-flow oracle in their “slow-trimming” procedure with an approximate one. We also avoid defining expanders and work exclusively with flow mixing, which simplifies the analysis.